OpenStudy (anonymous):

Write the quotient in standard form. 8-7i / 1-2i

OpenStudy (wasiqss):

multiply and divide by 1+2i , u ll get the answer

OpenStudy (anonymous):

\[)(1*2i)(1+2i)=1^2+2^2=5\] which is your denominator. the numerator will be \[(8-7i)(1+2i)\]